# Follow the Mass - A Concordance Picture of Tidal Disruption Events

### v1

@@ 4.3 Consensus structural results
Although the great majority of the bound debris stays relatively far from the black hole, small amounts can find their way to much closer. Again, the three extant simulations give similar results: at $t\simeq 1.4t_{0}$, both Ryu et al. (2023b) and Steinberg & Stone (2024) find $\sim 2-5\times 10^{-3}M_{\star}$ at radii $\lesssim r_{\rm p}/2$. Both numbers, as well as the fate of this matter, are uncertain due to the inner cut-offs employed by these simulations (see the next subsection for details). Although Price et al. (2024) do not quote a figure for the mass close to the black hole at $t\sim t_{0}$, their Figure 2 shows only a very small fraction of the total mass this close even at $t\simeq 9t_{0}$.
@@ 4.4 Limitations
To close this section, we acknowledge two limitations to the guidance we can derive from the existing global simulations. First, two of them impose unrealistic dynamics near the black hole. Ryu et al. (2023b) place an outflow boundary condition at a spherical radius $40r_{\rm g}$ from the black hole’s center. This means they cannot say what happens to matter passing through that surface, and they likely overestimate the net rate at which matter crosses the boundary. Rather than place a sharp boundary around the black hole, Steinberg & Stone (2024) instead force the gravitational potential to increase linearly with radius from the black hole out to $30r_{\rm g}$, where it switches to a Paczynski-Wiita form. The impact of this policy is to make the dynamics of gas close to the black hole unphysical: fluid elements have the wrong velocity and follow incorrect orbits. The motivation for both policies is to avoid excessively short timesteps. Like Ryu et al. (2023b), Price et al. (2024) employ general relativistic hydrodynamics (except for the star’s self-gravity), but differ by not requiring any central cut-out. Thus, this simulation may be more reliable than the other two for matter passing close to the black hole.
@@ 5.3 Optical depth and the cooling time
Because radiation transport in the slow-cooling regime is time-dependent, simple estimates of the luminosity, generally based on time-steady transfer, are subject to significant uncertainty. However, in this context, a simple estimate of the cooling time of individual fluid elements combined with the adiabaticity of the slow-cooling regime suggests that the luminosity of a given fluid element is held approximately constant as it moves around an orbit (Svirski et al., 2017). Estimated in this way, the luminosity released on a radial scale $r$ is $L(r)\sim P_{\rm rad}r^{2}c/\tau$. The radiation pressure varies $\propto r^{-4}$ when the volume of a moving fluid element is $\propto r^{3}$, as it is here due to the rough constancy of $h/r$, while $\tau$ for a specific fluid element is $\propto r^{-2}$. It follows that $L(r)$ is approximately independent of $r$. Additionally, as we will show in the next subsection, the heating rate in TDEs links this luminosity to the Eddington luminosity.
@@ 5.4.1 Luminosity
This much lower luminosity is greater than Eddington only for black holes with masses $\lesssim 5\times 10^{5}m_{\star}^{0.38}M_{\odot}$, and even then it exceeds Eddington by much less than the expectation based on relativistic efficiency.
Figure 2: Peak radiated luminosity as a function of $M_{\rm BH}$ and $m_{\star}$ on a logarithmic scale. We assume $h/r=0.3$, following the $t=t_{0}$ curve in Fig. 10d of Ryu et al. (2024). (Left) In units of $10^{44}$ erg s-1. (Right) In units of $L_{E}$.
To translate the rate at which energy is dissipated into a luminosity requires consideration of both the cooling time (see eqn. 14) and the duration of the heating, which is $\sim t_{0}$ for the bulk of the bound debris during the flare. The radiated luminosity $L_{\rm rad}=L_{\rm diss}$ when the cooling time is shorter than the time, i.e., $t_{0}$, in which the heat is dissipated. On the other hand, when $t_{\rm cool}>t_{0}$, the radiated luminosity is reduced by a factor $\sim t_{0}/t_{\rm cool}$ relative to the heating rate $L_{\rm diss}$. The peak luminosity (i.e., at $t=t_{0}$, when Ryu et al. (2024) found the cylindrical radius of the photosphere is at $r\approx a_{0}$) can then be written in a combined form that approximates both the slow and fast cooling regimes:
@@
A small fraction of optical/UV TDEs shows still higher luminosities, reaching as high as $5\times 10^{45}$ erg s-1 (Hammerstein et al., 2023). As explained in Sec. 5.8, very luminous events are overrepresented in a magnitude-limited sample, so their real fraction is significantly smaller than the fraction observed. We speculate that these are the rare TDEs in which the star plunges to $r_{\rm p}\lesssim 10r_{\rm g}$, a regime in which general relativistic apsidal precession is substantial and changes the character of TDEs, likely yielding higher luminosity (see e.g. Krolik et al., 2020).
@@ 5.4.3 Duration
In the rapid-cooling regime, the optical/UV lightcurve should track the rate at which the debris’ internal energy would increase in the absence of cooling. This rate is driven by the mass return rate, but its decline following the peak in the mass return rate is more gradual, as fluid elements continue to suffer shocks as they orbit around the black hole. Because the placement and strength of these shocks depend on details of the flow, the degree to which the decline is slower than the decline in the mass return rate could vary from case to case. Nonetheless, in this regime the duration of the flare should generally be a few $t_{0}$.
@@ 5.5.4 Duration
As discussed in Sec. 3.2 and originally pointed out in Ryu et al. (2020c) (see its Fig. 8a), the increase of $\Xi$ with $M_{\star}$ causes $t_{0}$ to be almost independent of $M_{\star}$. For this reason, the timescale of a flare is a very weak indicator of $M_{\star}$. It is, on the other hand, $\propto M_{{}_{\rm BH}}^{0.6}$. Unfortunately, the persistence of shock heating for several $t_{0}$ means that the period of greatest luminosity is a few $t_{0}$. Quantitative prediction of the flare duration therefore demands a better calibration of “a few" than the simulations to date can give.
@@ 5.6 Late-time evolution
Over longer timescales, years or more, nearly all the bound mass should eventually accrete. If most accretes with high radiative efficiency, the total energy emitted over long timescales could be as large as $\sim 10^{53}m_{\star}$ erg, several hundred times the energy emitted during the flare. However, this could take a very long time. On the other hand, the radiated energy is reduced to the extent that some of the gas has especially low angular momentum matter; such gas can plunge ballistically, passing through the event horizon without significant dissipation (Svirski et al., 2017).
@@
If inflow is driven by angular momentum transport, it requires internal stresses, which are generally due to correlated MHD turbulence. Building the turbulence from scratch takes $\sim 5-10$ orbital periods, i.e., $\sim 5-10t_{0}$, whether the orbits are circular or elliptical (Stone et al., 1996; Hawley & Krolik, 2001; Chan et al., 2023). Even after the turbulence reaches nonlinear saturation, the characteristic inflow time is $\sim 10(r/h)^{2}$ orbital periods, where $h/r$ is the disk aspect ratio. Thus, one might expect (Shiokawa et al., 2015) that the timescale on which most of the debris accretes is $t_{\rm late}\gtrsim 2\times 10^{3}(h/0.1r)^{-2}\,t_{0}$, which translates to
$t_{\rm late}\sim 200~{}\Xi^{-3/2}(h/0.1r)^{-2}m_{\rm BH,6}^{1/2}m_{\star}^{0.82}\hbox{~{}yr}$ (20)
and luminosity (if the accretion is radiatively efficient)
[…]
Here our fiducial $h/r=0.1$ rather than 0.5 because cooling should lead to a diminishing scale height. Even though the MHD-mediated inflow is slow at the large distance where most of the mass initially resides, there is so much more of it than the mass that went quickly into a small disk that the luminosity generated at late times is down by only a single order of magnitude from the peak:
$\frac{L_{\rm late}}{L_{\rm rad,peak}}\approx 0.1\Xi(h/0.1r)^{2}m_{{}_{\rm BH,6}}^{-1/3}m_{\star}^{-0.26}.$ (22)
Moreover, this ratio is only weakly dependent on $M_{\star}$ and $M_{{}_{\rm BH}}$, and $L_{\rm late}$ should also be rather slowly-varying because it takes many years to deplete the mass reservoir on the $r\sim a_{0}$ scale. However, there may be another source of time-dependence. As shown by Chan et al. (2023), outward angular momentum transport in eccentric disks can accelerate accretion by making their inner regions even more eccentric. This effect might tilt the late-time lightcurve, raising the luminosity at the start of MHD-driven accretion and diminishing it somewhat at still later times. It should also be noted that once MHD-mediated accretion begins (at $t\gtrsim 5-10t_{0}$), the hydrodynamical mechanisms studied by Price et al. (2024) may not be good predictors of the inflow rate.
Thus, the optical/UV lightcurve can be expected to flatten out on roughly year-timescales, plateauing at a level 0.1$\sim$ 0.1$\times$ 0.1 × the peak (for rapidly-cooling events). In fact, the lightcurves of many TDEs do flatten out at times $\sim 1$ yr past peak, and when they do, the most common value of $L_{\rm late}/L_{\rm rad,peak}$ is $\sim 0.1$ (van Velzen et al., 2019b; Hammerstein et al., 2023; Yao et al., 2023; Mummery et al., 2024).
It is also important to note that the total energy radiated at late times, $\sim 3\times 10^{50}\Xi^{3/2}(h/0.1r)^{2}m_{{}_{\rm BH,6}}^{-1/2}m_{\star}^{0.18}(\Delta t/1$ yr) erg, is comparable to the total energy radiated during the flare. This fact, in itself, demonstrates that the majority of the debris mass cannot have been brought to smaller radii during the flare.
As a final remark, we emphasize that the initial deposition of nearly all the debris mass on the $\sim a_{0}$ scale implies directly that the accretion flow in TDEs is very far from inflow equilibrium—in other words, the accretion rate varies substantially with distance from the black hole—everywhere except possibly very near the black hole (see Sec. 5.7).
@@ 5.7.1 Luminosity
Taken at face value, $L_{\rm defl}$ declines with $M_{\rm BH}$ and increases with $M_{\star}$, and the explicit mass-scalings for both are steepened by the trends in $\Xi$. However, there are numerous caveats and uncertainties attached to this estimate. The mass deflection rate itself remains poorly determined, and its dependence on $m_{\star}$ and $m_{{}_{\rm BH,6}}$ is unknown. In addition, there are several ways the radiative efficiency of the deflected gas could be less than the canonical relativistic efficiency of 0.1. As mentioned briefly at the end of Sec. 4, the debris orbits have energy much greater than that of a circular orbit with their angular momentum. To join a small, circular orbit therefore requires the loss of a large amount of energy, and the simulations do not exhibit a mechanism for this. If the debris orbits lose a small amount of their angular momentum, the associated gas can plunge directly across the event horizon without having lost any of its kinetic energy to dissipation (Svirski et al., 2017). Even if the angular momentum of the deflected debris is large enough for it to go into orbit around the black hole, it may still be small enough that Reynolds stresses suffice to push the gas onto a plunging orbit; there may be associated dissipation, but too little time to radiate (Avara et al., 2023). Even if there is dissipation and photons are emitted, high optical depth may suppress photon escape (Price et al., 2024) or reprocess the radiation, as we will discuss in the next subsection.
@@ 5.7.2 Spectrum
Note that in this sense, reprocessing of photons radiated closer to the black hole can be important. However, it acts upon a far smaller luminosity than originally envisioned. This sort of reprocessing also contrasts with other reprocessing scenarios in that here the gas doing the reprocessing is the bulk of the bound debris, which remains at distances from the black hole $\sim 10^{3}r_{\bf g}$, rather than the debris made unbound at the time of disruption (Strubbe & Quataert, 2009) or a wind driven by super-Eddington radiation (Strubbe & Quataert, 2011; Metzger & Stone, 2016).

### v2

@@ 4.3 Consensus structural results
Although the great majority of the bound debris stays relatively far from the black hole, small amounts can find their way much closer. Again, the three extant simulations give similar results: at $t\simeq 1.4t_{0}$, both Ryu et al. (2023b) and Steinberg & Stone (2024) find $\sim 2-5\times 10^{-3}M_{\star}$ at radii $\lesssim r_{\rm p}/2$. Both numbers, as well as the fate of this matter, are uncertain due to the inner cut-offs employed by these simulations (see the next subsection for details). Although Price et al. (2024) do not quote a figure for the mass close to the black hole at $t\sim t_{0}$, their Figure 2 shows only a very small fraction of the total mass this close even at $t\simeq 9t_{0}$.
@@ 4.4 Limitations
To close this section, we acknowledge three limitations to the guidance we can derive from the existing global simulations. First, two of them impose unrealistic dynamics near the black hole. Ryu et al. (2023b) place an outflow boundary condition at a spherical radius $40r_{\rm g}$ from the black hole’s center. This means they cannot say what happens to matter passing through that surface, and they likely overestimate the net rate at which matter crosses the boundary. Rather than place a sharp boundary around the black hole, Steinberg & Stone (2024) instead force the gravitational potential to increase linearly with radius from the black hole out to $30r_{\rm g}$, where it switches to a Paczynski-Wiita form. The impact of this policy is to make the dynamics of gas close to the black hole unphysical: fluid elements have the wrong velocity and follow incorrect orbits. The motivation for both policies is to avoid excessively short timesteps. Like Ryu et al. (2023b), Price et al. (2024) employ general relativistic hydrodynamics (except for the star’s self-gravity), but differ by not requiring any central cut-out. Thus, this simulation may be more reliable than the other two for matter passing close to the black hole.
@@
The third limitation is that all three simulations considered events with $r_{\rm p}>10r_{\rm g}$. When the pericenter is smaller, relativistic apsidal precession in a single pericenter passage is a radian or more, so that strong shocks take place close to the pericenter scale, and a great deal more heat is dissipated than when the apsidal precession is weaker and the shocks take place much farther away. This contrast can qualitatively change the character of the event, especially when $r\lesssim 6r_{\rm g}$ (Ryu et al., 2023a). Pericenter passages closer than $\sim 10r_{\rm g}$ occur in only a small fraction of disruptions by black holes with mass $\lesssim 10^{7}M_{\odot}$, but in nearly all flare-producing events with $M_{\rm BH}\gtrsim 10^{7}M_{\odot}$ (Ryu et al., 2020c; Krolik et al., 2020). For this reason, the consensus results we focus on here may need significant revision when $M_{\rm BH}\gtrsim 10^{7}M_{\odot}$, and our statements about scaling relations may fail for $M_{\rm BH}\gtrsim 10^{7}M_{\odot}$.
@@ 5.3 Optical depth and the cooling time
Because radiation transport in the slow-cooling regime is time-dependent, simple estimates of the luminosity, generally based on time-steady transfer, are subject to significant uncertainty. However, in this context, a simple estimate of the cooling time of individual fluid elements combined with the adiabaticity of the slow-cooling regime suggests that the luminosity of a given fluid element is held approximately constant as it moves around an orbit (Svirski et al., 2017). Estimated in this way, the luminosity released on a radial scale $r$ is $L(r)\sim P_{\rm rad}r^{2}c/\tau$, where $P_{\rm rad}$ is the radiation pressure. With $P_{\rm rad}\propto r^{-4}$ when the volume of a moving fluid element is $\propto r^{3}$, as it is here due to the rough constancy of $h/r$, for a specific fluid element $\tau\propto r^{-2}$. It follows that $L(r)$ is approximately independent of $r$. Additionally, as we will show in the next subsection, the heating rate in TDEs links this luminosity to the Eddington luminosity.
@@ 5.4.1 Luminosity
This much lower heating rate is greater than Eddington only for black holes with masses $\lesssim 5\times 10^{5}m_{\star}^{0.38}M_{\odot}$, and even then it exceeds Eddington by much less than the expectation based on relativistic efficiency.
Figure 2: Peak radiated luminosity as a function of $M_{\rm BH,6}$ and $m_{\star}$ on a logarithmic scale ranging from 0.1 to 10 for both quantities. We assume $h/r=0.3$, following the $t=t_{0}$ curve in Fig. 10d of Ryu et al. (2024). (Left) In units of $10^{44}$ erg s-1. (Right) In units of $L_{E}$.
To translate the rate at which energy is dissipated into a radiated luminosity requires consideration of both the cooling time (see eqn. 14) and the duration of the heating, which is $\sim t_{0}$ for the bulk of the bound debris during the flare. The radiated luminosity $L_{\rm rad}=L_{\rm diss}$ when the cooling time is shorter than the time, i.e., $t_{0}$, in which the heat is dissipated. On the other hand, when $t_{\rm cool}>t_{0}$, the radiated luminosity is reduced by a factor $\sim t_{0}/t_{\rm cool}$ relative to the heating rate $L_{\rm diss}$. The peak luminosity (i.e., at $t=t_{0}$, when Ryu et al. (2024) found the cylindrical radius of the photosphere is at $r\approx a_{0}$) can then be written in a combined form that approximates both the slow and fast cooling regimes:
@@
More importantly, for $\sim 0.3\lesssim m_{\star}\lesssim 3$, they are closely consistent with the peak luminosities exhibited by the bulk of the observed cases (see Sec. 2 and the references cited there). Such good agreement with the observed peak luminosities gives very strong support to this picture because its prediction of the luminosity scale has no free parameters or guessed initial conditions for the debris; it is based entirely on physical calculations beginning with a main-sequence star passing a black hole. On the other hand, a small fraction of optical/UV TDEs shows still higher luminosities, reaching as high as $5\times 10^{45}$ erg s-1 (Hammerstein et al., 2023a). As explained in Sec. 5.8, very luminous events are overrepresented in a magnitude-limited sample, so their real fraction is significantly smaller than the fraction observed. Because the black hole masses associated with these very luminous events tend to be $\gtrsim 10^{7}M_{\odot}$ (Hammerstein et al., 2023b; Mummery et al., 2024; Yao et al., 2023), we speculate that these are the rare TDEs in which the star plunges to $r_{\rm p}\lesssim 10r_{\rm g}$, the large apsidal precession regime described at the end of Sec. 4.4. The different dynamics of that regime may lead to their higher luminosity.
@@ 5.4.3 Duration
In the rapid-cooling regime, the optical/UV lightcurve should track the rate at which the debris’ internal energy would increase in the absence of cooling. This rate is driven by the mass return rate, but its decline following the peak in the mass return rate is more gradual, as fluid elements continue to suffer shocks as they orbit around the black hole. Because the placement and strength of these shocks depend on details of the flow, the degree to which the decline is slower than the decline in the mass return rate could vary from case to case, which means a small set of simulations is insufficient to predict the specific shape of the declining lightcurve. Nonetheless, in this regime the duration of the flare should generally be a few $t_{0}$.
@@ 5.5.4 Duration
As discussed in Sec. 3.2 and originally pointed out in Ryu et al. (2020d) (see its Fig. 8a), the increase of $\Xi$ with $M_{\star}$ causes $t_{0}$ to be almost independent of $M_{\star}$. For this reason, the timescale of a flare is a very weak indicator of $M_{\star}$. It is, on the other hand, $\propto M_{{}_{\rm BH}}^{0.6}$. Unfortunately, the persistence of shock heating for several $t_{0}$ means that the period of greatest luminosity is a few $t_{0}$. Quantitative prediction of the flare duration therefore demands a better calibration of “a few" than the simulations to date can give.
@@ 5.6 Late-time evolution
Over longer timescales, years or more, nearly all the bound mass should eventually accrete. If most accretes with high radiative efficiency, the total energy emitted over long timescales could be as large as $\sim 10^{53}m_{\star}$ erg, several hundred times the energy emitted during the flare. Indeed, even observing TDEs for a few years past the flare has shown that the total energy radiated at late times, $\sim 3\times 10^{50}\Xi^{3/2}(h/0.1r)^{2}m_{{}_{\rm BH,6}}^{-1/2}m_{\star}^{0.18}(\Delta t/1$ yr) erg, is comparable to the total energy radiated during the flare. This fact, in itself, confirms what the three global simulations have shown: that the majority of the debris mass cannot have been brought to smaller radii during the flare, and that the inward progress of the debris is generically very slow. This conclusion is contrary to the assumption made in many models that the matter is deposited at small radius ($r\sim r_{\rm p}$) early on and then spreads both inward and outward (Cannizzo et al., 1990; Shen & Matzner, 2014; Mummery & Balbus, 2020).
Much about late-time behavior is beyond what can be predicted from the three simulations’ results. For example, whether the entire $\sim 10^{53}m_{\star}$ erg is eventually radiated is unclear. Among numerous uncertainties, the radiated energy could be reduced to the extent that some of the gas acquires especially low angular momentum; such gas can plunge ballistically, passing through the event horizon without significant dissipation (Svirski et al., 2017).
@@
On the other hand, certain properties of the late-time luminosity can be discerned from what we now know. One robust characteristic is that the inflow time is long because the mass is deposited far from the black hole. If inflow is driven by angular momentum transport, it requires internal stresses, which are generally due to correlated MHD turbulence. Building the turbulence from scratch takes $\sim 5-10$ orbital periods, i.e., $\sim 5-10t_{0}$, whether the orbits are circular or elliptical (Stone et al., 1996; Hawley & Krolik, 2001; Chan et al., 2023). Even after the turbulence reaches nonlinear saturation, there is considerable uncertainty about the magnitude of the resulting stresses. The inflow time is conventionally parameterized in terms of a pair of ratios, neither one easy to predict: $t_{\rm inflow}\sim\alpha^{-1}(r/h)^{2}$ orbital periods, where $\alpha$ is the time- and azimuthally-averaged ratio of the vertically-integrated stress to the vertically-integrated pressure. Guessing at these parameters, one might expect (Shiokawa et al., 2015) that the timescale on which most of the debris accretes is $t_{\rm late}\gtrsim 2\times 10^{3}(\alpha/0.1)^{-1}(h/0.1r)^{-2}\,t_{0}$. Here we scale $h/r$ in terms of 0.1 because the inflow timescale is likely longer than the cooling time. In terms of TDE parameters, this expression for the inflow time translates to
$t_{\rm late}\sim 200~{}\Xi^{-3/2}(\alpha/0.1)^{-1}(h/0.1r)^{-2}m_{\rm BH,6}^{1/2}m_{\star}^{0.82}\hbox{~{}yr}.$ (20)
We emphasize that the fiducial quantities in this expression are plausible, but hardly well-determined.
If, as one would expect, $t_{\rm late}$ is at least several years, it immediately follows that the luminosity radiated as the matter accretes toward the black hole should not have major trends on timescales shorter than several years. In other words, the late-time luminosity should flatten out. In fact, the lightcurves of many TDEs do just this at times $\sim 1$ yr past peak. When they do, the most common value of $L_{\rm late}/L_{\rm rad,peak}$ measured in terms of $\nu L_{\nu}$ in the NUV is $\sim 0.1$, but in some cases it is smaller by a factor of several to ten (van Velzen et al., 2019b; Yao et al., 2023; Mummery et al., 2024). That the luminosity remains within one or two orders of magnitude of the flare peak for many years can be readily understood. The radiative efficiency of matter accreting onto a black hole is $\sim 10^{2}\times$ greater than the radiative efficiency of shocks on the $r\sim a_{0}$ scale, and this advantage comes close to compensating for the disadvantage of a timescale $\sim 10^{3}\times$ longer.
Substantial scatter in the ratio of late-time to flare amplitude might be expected from a number of causes. The saturation amplitude of the MHD turbulence may depend on accretion conditions. With much of the debris remaining at distances $\gtrsim 10^{3}r_{\rm g}$ while the accretion luminosity is produced, as usual, close to the black hole, some fraction of solid angle may be blocked by the debris. The debris orbits at large distance from the black hole should align with the stellar orbit, but Lense-Thirring torques acting closer to the black hole may modulate the inflow or lead to dissipation due to mechanisms other than damping of MHD turbulence. This list is undoubtedly incomplete.
@@ 5.7.1 Luminosity
Taken at face value, $L_{\rm defl}$ declines with $M_{\rm BH}$ and increases with $M_{\star}$, and the explicit mass-scalings for both are steepened by the trends in $\Xi$. However, there are numerous caveats and uncertainties attached to this estimate. The mass deflection rate itself remains poorly determined, and its dependence on $m_{\star}$ and $m_{{}_{\rm BH,6}}$ is unknown. In addition, there are several ways the radiative efficiency of the deflected gas could be less than the canonical relativistic efficiency of 0.1. As mentioned briefly at the end of Sec. 4, the debris orbits have energy much greater than that of a circular orbit with their angular momentum. To join a small, circular orbit therefore requires the loss of a large amount of energy, and the simulations do not exhibit a mechanism for this. If the debris orbits lose a small amount of their angular momentum, the associated gas can plunge directly across the event horizon without having lost any of its kinetic energy to dissipation (Svirski et al., 2017). Even if the angular momentum of the deflected debris is large enough for it to go into orbit around the black hole, it may still be small enough that Reynolds stresses suffice to push the gas onto a plunging orbit; there may be associated dissipation, but too little time to radiate (Chan et al., 2023). Even if there is dissipation and photons are emitted, high optical depth may suppress photon escape (Price et al., 2024) or reprocess the radiation, as we will discuss in the next subsection.
@@ 5.7.2 Spectrum
Note that in this sense, reprocessing of photons radiated closer to the black hole can be important. However, it acts upon a far smaller luminosity than originally envisioned. This sort of reprocessing also contrasts with other reprocessing scenarios in that here the gas doing the reprocessing is the bulk of the bound debris, which remains at distances from the black hole $\sim 10^{3}r_{\rm g}$, rather than the debris made unbound at the time of disruption (Strubbe & Quataert, 2009) or a wind driven by super-Eddington radiation (Strubbe & Quataert, 2011; Metzger & Stone, 2016).